OpenStudy (ray_sihota):

A boy scout troop is selling Christmas trees to raise money for a trip. On the First day the they sold 79 trees. The balsam firs sold for $23 and Norwegian pins were $28. The receipts for the day totaled$1942. How many of each trees of each type was sold.

OpenStudy (anonymous):

@superhelp101

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

maybe these people can help you.

OpenStudy (ray_sihota):

thanks

OpenStudy (anonymous):

Welcome... :)

OpenStudy (imstuck):

You need to have a system of equations to solve this. First relate the NUMBER of trees sold to each other, and then for the second equation, relate the MONEY to each other.

OpenStudy (imstuck):

Call the Balsam Firs B and the Norweigan pines N, ok? And they sold a total of 79 trees. So our first equation, relating the NUMBER of trees sold to the total is B + N = 79

OpenStudy (imstuck):

That means that the total number of Balsams sold PLUS the total number of Norweigians sold is 79.

OpenStudy (imstuck):

Now let's do the money part.

OpenStudy (ray_sihota):

money part will be 23+ 28=1942

OpenStudy (imstuck):

Each Balsam cost $23 and each Norwegian cost $28, and the total money brought in was $1942. So for our Balsam, we have the number of Balsams, which we don't know, we call those B, times how much per Balsam which is $23. So for a Balsam, we have 23B. Add that to the cost of a Norwegian, which is 28 per Norwegian. So Norwwegians, money-wise, is 28N. The total sold is 1942. So the money part is 23B + 28N = 1942

OpenStudy (imstuck):

Our system of equations is this now: B + N = 79 23B + 28N = 1942

OpenStudy (ray_sihota):

Ah got it thank so much

OpenStudy (imstuck):

Did you solve it?

OpenStudy (ray_sihota):

I thought i did but i am stuck then do I just use the addition method to solve it

OpenStudy (ray_sihota):

yes -23(b+w=79) = -23b-23n-1817+ 23B+28n=1942= 5n=125 125/5=25 w=25 b+25=79 -25 both sides b=54